a) In a circuit containing inductance 𝐿, resistance 𝑅 and voltage 𝐸, the current 𝐼 is


given by 𝐸 = 𝑅 𝐼 + 𝐿


𝑑𝐼


𝑑𝑑


. Given 𝐿 = 640 β„Ž , 𝑅 = 250 𝛺 and 𝐸 =


500 π‘£π‘œπ‘™π‘‘ . 𝐼 being 0 when 𝑑 = 0. Find the time 𝑑 that elapses, before the current


𝐼 reaches 90% of its maximum value.


[5]


b) Solve the system:


𝑑π‘₯


𝑑𝑑


+ π‘₯ βˆ’ 𝑦 = 𝑑𝑒


𝑑


, 2𝑦 βˆ’


𝑑π‘₯


𝑑𝑑


+


𝑑𝑦


𝑑𝑑


= 𝑒

1
Expert's answer
2021-12-30T02:53:24-0500

a) Differential equation of a LR circuitΒ 


"L\dfrac{dI}{dt}+RI=E""\dfrac{dI}{dt}+\dfrac{R}{L}I=\dfrac{E}{L}"

Integrating factor


"\mu(t)=e^{\int(R/L)dt}=e^{(R/L)t}"

"e^{(R/L)t}(\dfrac{dI}{dt}+\dfrac{R}{L}I)=e^{(R/L)t}\dfrac{E}{L}"

"d(e^{(R/L)t}I)=e^{(R/L)t}\dfrac{E}{L}dt"

Integrate


"\int d(e^{(R/L)t}I) =\int e^{(R/L)t}\dfrac{E}{L}dt"

"e^{(R/L)t}I=\dfrac{E}{R}e^{(R/L)t}+C"

"I(t)=\dfrac{E}{R}+Ce^{-(R/L)t}"

"I(0)=0"

"0=\dfrac{E}{R}+C"

"C=-\dfrac{E}{R}"

"I(t)=\dfrac{E}{R}(1-e^{-(R/L)t})"

Given "E=500V, R=250\ Ohm, L=640\ H"

Substitute


"I(t)=\dfrac{500}{250}(1-e^{-(250/640)t})"




"I(t)=2(1-e^{-(25\/64)t})"

Find the time 𝑑 that elapses, before the current "I" reaches 90% of its maximum value


"I(t_{1})=2(1-e^{-(25\/64)t_1})=2(0.9)"

"e^{-(25\/64)t_1}=0.1"

"\dfrac{25}{64}t_1=\ln (10)"

"t_1=2.56\ln(10) \ s"

"t_1\approx5.8946\ s"

b)


"\dfrac{dx}{dt}+x-y=te^t"

"2x-\dfrac{dx}{dt}+\dfrac{dy}{dt}=e^t"

Differentiate the first equation with respect to "t"


"\dfrac{d^2x}{dt^2}+\dfrac{dx}{dt}-\dfrac{dy}{dt}=te^t+e^t"

Substitute


"\dfrac{dy}{dt}=\dfrac{dx}{dt}-2x+e^t"

"\dfrac{d^2x}{dt^2}+\dfrac{dx}{dt}-\dfrac{dx}{dt}+2x-e^t=te^t+e^t"

"\dfrac{d^2x}{dt^2}+2x=te^t+2e^t"

Corresponding homogeneous differential equation


"\dfrac{d^2x}{dt^2}+2x=0"

Auxiliary equation


"r^2+2=0"

"r=\pm\sqrt{2}i"

The general solution of the homogeneous differential equation is


"x_h=c_1\cos(\sqrt{2}t)+c_2\sin(\sqrt{2}t)"

Find the partial solution of the nonhomogeneous differential equation


"\tilde{x}=Ate^t+Be^t"

"\tilde{x}'=Ate^t+Ae^t+Be^t"

"\tilde{x}''=Ate^t+2Ae^t+Be^t"

Substitute


"Ate^t+2Ae^t+Be^t+2Ate^t+2Be^t"

"=te^t+2e^t"

"A=\dfrac{1}{3}, B=\dfrac{4a}{9}"

The general solution of the nonhomogeneous differential equation is


"x(t)=c_1\cos(\sqrt{2}t)+c_2\sin(\sqrt{2}t)+\dfrac{1}{3}te^t +\dfrac{4}{9}e^t""\dfrac{dx}{dt}=-c_1\sqrt{2}\sin(\sqrt{2}t)+c_2\sqrt{2}\cos(\sqrt{2}t)""+\dfrac{1}{3}te^t +\dfrac{1}{3}e^t +\dfrac{4}{9}e^t"

"y= \dfrac{dx}{dt}+x-te^t"

"y=-c_1\sqrt{2}\sin(\sqrt{2}t)+c_2\sqrt{2}\cos(\sqrt{2}t)"

"+\dfrac{1}{3}te^t +\dfrac{7}{9}e^t+c_1\cos(\sqrt{2}t)+c_2\sin(\sqrt{2}t)"

"+\dfrac{1}{3}te^t +\dfrac{4}{9}e^t-te^t"


"x(t)=c_1\cos(\sqrt{2}t)+c_2\sin(\sqrt{2}t)+\dfrac{1}{3}te^t +\dfrac{4}{9}e^t"

"y(t)=(c_1+c_2\sqrt{2})\cos(\sqrt{2}t)+(-c_1\sqrt{2}+c_2)\sin(\sqrt{2}t)"

"-\dfrac{1}{3}te^t +\dfrac{11}{9}e^t"